Solve the following ordinary differential equations

(a) dy/dx+ycotx=power of e is cosx
For x=π/2 y=-2

(b) d square y/dx square + dy/dx + y =0
1
Expert's answer
2020-02-24T10:47:22-0500

a)


"\frac{dy}{dx}+y\cot x=e^{cosx}"

"p=\cot x, q=e^{cosx}"

"IF=\exp\int pdx=\exp\int \cot xdx=\exp (\ln\sin x)"

"IF=\sin x"

"y\sin x=\int e^{cosx}\sin x dx=c-e^{cosx}"

"-2\sin 90=c-e^{cos90}\to c=-1"

"y=-\frac{1+e^{cos90}}{\sin x}"

b)


"\frac{d^2y}{dx^2}+\frac{dy}{dx}+y=0\to y=e^{\lambda x}"

"\lambda ^2+\lambda +1=0"

"\lambda_{1,2}=\frac{-1\pm \sqrt{3}}{2}"

"y=c_1e^{-\frac{x}{2}}\sin\left (\frac{\sqrt{3}}{2}x\right)+c_2e^{-\frac{x}{2}}\cos\left (\frac{\sqrt{3}}{2}x\right)"